OpenStudy (anonymous):

12.03 × 0.76= 7218+84210= 914.28 What did I do wrong, and how can I correct it?

OpenStudy (anonymous):

lol what is this?

OpenStudy (anonymous):

My homework

OpenStudy (uri):

@kms13 What's the question?

OpenStudy (anonymous):

Plz help me

OpenStudy (wolf1728):

I think I have an answer to the unasked question. A) 12.03 + .76 = 9.1428 B) 7218 + 84210 = 91428 9.1428 is lower than 914.28 by a factor of .01 91428 is larger than 914.28 by a factor of 100

OpenStudy (wolf1728):

oops A) should be 12.03 * .76

OpenStudy (jdoe0001):

beware of the clairvoyant abilities of wolf1728

OpenStudy (jdoe0001):

:P

OpenStudy (wolf1728):

So, if I have perceived the spirit world correctly: 12 * 0.76 x 100 = 914.28 (7218 + 84210) / 100 = 914.28